# DM-Codec: Distilling Multimodal Representations for Speech Tokenization

### v1

@@ 2.2 Model Details
Our framework builds upon the Residual Vector Quantizer with Generative Adversarial Networks (RVQ-GAN) architecture, incorporating state-of-the-art components and novel distillation techniques. The core of our model consists of an Encoder $\mathbf{E}$ and Decoder $\mathbf{D}$ with an RVQ architecture, inspired by Encodec (Défossez et al., 2022) and SpeechTokenizer (Zhang et al., 2024a). Moreover, we employ a multi-discriminator framework, comprising: Multi-Scale Discriminator (MSD), Multi-Period Discriminator (MPD), and Multi-Scale Short-Time Fourier Transform (MS-STFT) Discriminator, adopted from HiFi-Codec (Yang et al., 2023) and HiFi-GAN (Kong et al., 2020). This foundation provides a robust basis for speech quantization. To further enhance the quantizer with distilled multimodal representations, we use wav2vec 2.0 (wav2vec2-base-960h) as $M_{STT}$ (Baevski et al., 2020), BERT (bert-base-uncased) as $M_{LM}$ (Devlin et al., 2019), and HuBERT (hubert-base-ls960) as $M_{SM}$ (Hsu et al., 2021). We extract the quantized output from the first layer of the RVQ (RVQ-1) for LM-guided distillation and the average of the quantized features across all eight layers (RVQ-1:8) for SM-guided distillation to calculate the distillation loss.
2.3 Model Components
Residual Vector Quantizers. The Residual Vector Quantizer (RVQ) plays a central role in our tokenization process, quantizing the encoder’s outputs. Our implementation is inspired by the training procedures described in Encodec (Défossez et al., 2022) and SpeechTokenizer (Zhang et al., 2024a). The RVQ projects input vectors to the most similar entry in a codebook, and the residual is calculated and processed in subsequent quantization steps, each utilizing a different codebook. The codebook entries are updated using an exponential moving average (EMA) with a decay rate of 0.99 for the matched item, while unmatched entries are replaced by candidates from the current batch. To ensure proper gradient flow during training, we employ a straight-through estimator. A commitment loss is also computed and added to the total training loss to promote stability. In our experiments, we utilize a codebook size of 1024 and 8 quantization levels.
2.4 Training Objective
@@
Our training strategy employs a GAN-guided framework, following methodologies established in recent work (Zhang et al., 2024a; Yang et al., 2023). In addition to the distillation loss described in Section 2.1, we utilize reconstruction losses, adversarial and feature matching losses, and a commitment loss to guide the learning process. For the original speech $\mathbf{x}$ and the reconstructed speech $\hat{\mathbf{x}}$, we calculate the losses as described below.
@@
Reconstruction Loss. To ensure that the model preserves the key attributes of speech, we employ both time-domain and frequency-domain reconstruction losses. The time-domain loss $\mathcal{L}_{t}$ is computed as the L1 distance between $\mathbf{x}$ and $\hat{\mathbf{x}}$. For the frequency-domain loss $\mathcal{L}_{f}$, we combine L1 and L2 losses over 64-bin Mel-spectrograms $\text{Mel}_{i}$, with varying window sizes of $2^{i}$, hop lengths of $42^{i}/$ 4, and scales $e=\{5,\ldots,11\}$.
@@
$\mathcal{L}_{f}$ $=$ $\sum_{i\in e}\left(\|\text{Mel}_{i}(\mathbf{x})-\text{Mel}_{i}(\hat{\mathbf{x}})\|_{1}+\|\text{Mel}_{i}(\mathbf{x})-\text{Mel}_{i}(\hat{\mathbf{x}})\|_{2}\right)$ (5)
Adversarial Loss. The adversarial loss promotes the generator to produce realistic and indistinguishable speech. We apply a hinge loss formulation to compute the adversarial loss for both the generator $\mathcal{L}_{g}$ and the discriminator $\mathcal{L}_{d}$. These losses are computed across all three discriminators: the multi-scale discriminator (MSD), multi-period discriminator (MPD), and the multi-scale STFT guided (MS-STFT) discriminator.
@@
$\mathcal{L}_{d}$ $=$ $\frac{1}{N}\sum_{n=1}^{N}\left(\max(1-R_{n}(\mathbf{x}),0)+\max(1+R_{n}(\hat{\mathbf{x}}),0)\right)$ (7)
@@
$\mathcal{L}_{G}=\lambda_{L/LS}\mathcal{L}_{L/LS}+\lambda_{t}\mathcal{L}_{t}+\lambda_{f}\mathcal{L}_{f}+\lambda_{g}\mathcal{L}_{g}+\lambda_{fm}\mathcal{L}_{fm}+\lambda_{w}\mathcal{L}_{w}$ (10)
@@
3 Experimental Setup
Dataset. We trained DM-Codec using the LibriSpeech training set of 100 hours of clean speech (Panayotov et al., 2015). This dataset was selected primarily because of its successful use for training and evaluation in various speech tokenizer and modeling tasks (Zhang et al., 2024a; Ju et al., 2024; Hsu et al., 2021). Before training, we made the data uniform by randomly cropping each sample to three seconds and ensuring a consistent sample rate of 16 Hz.
Training. We trained DM-Codec utilizing 2 to 4 A100 GPUs until the model converged within 100 epochs. The batch size ranged from 6 to 20, depending on GPU resource availability. We applied a learning rate of $1\times 10^{-4}$ using the Adam optimizer with a 0.98 learning rate decay. The embedding size was set to 1024 for RVQ and 768 for the LM and SM. For all experiments, we used a random seed of 42 to ensure reproducibility. We also share our training code with the entire configuration file and a docker file to reproduce the training environment, which can be accessed through the GitHub link referenced in the abstract.
Baselines. We compared DM-Codec with the baseline speech tokenizers: EnCodec (Défossez et al., 2022), SpeechTokenizer (Zhang et al., 2024a), and FACodec (NaturalSpeech3) (Ju et al., 2024). We reproduced SpeechTokenizer using the official training code and used official model checkpoints of EnCodec (EnCodec_24khz_6kpbs) and FACodec as the baselines.
Evaluation Dataset. To evaluate DM-Codec, we randomly selected 300 audio samples from the LibriSpeech test subset, following a similar practice of sampling test data used in our baselines (Zhang et al., 2024a; Zeghidour et al., 2021) and to align the experimental setup with that of SpeechTokenizer. In our experiments, we sampled the test subset of LibriSpeech using a random seed of 42. We also evaluated the baseline models with the same sampled test dataset for a fair comparison.
Evaluation Metrics. To evaluate DM-Codec, we employed different metrics suited to get insights into various aspects of information and quality preservation in the reconstructed speech. First, we used the Word Error Rate (WER) and Word Information Lost (WIL) metrics to evaluate context preservation by calculating the amount of word-level transcription errors and key information missing in transcription, respectively. For these metrics, we used the Whisper (whisper-medium) (Radford et al., 2023) model to extract the transcription from the reconstructed speech. To provide a fairer comparison and indicate the level of transcription error by the Whisper model, we also included the Groundtruth WER and WIL scores for the Whisper’s transcribed text from the original speech versus the true text. Next, we assessed the acoustic and semantic information preservation using the ViSQOL (Virtual Speech Quality Objective Listener) (Hines et al., 2012) and Short-Time Objective Intelligibility (STOI) metrics, respectively. The ViSQOL metric measures the similarity between a reference and a test speech sample using a spectro-temporal measure and produces a MOS-LQO (Mean Opinion Score - Listening Quality Objective) score ranging from 1 (worst) to 5 (best). For this metric, we used the wideband model suited for speech evaluation. Lastly, the STOI metric evaluates the perceived intelligibility of speech by analyzing short-time correlations between original and reconstructed speech, with scores ranging from 0 to 1.
@@ 3 Experimental Setup
Tokenizer WER $\downarrow$ WIL $\downarrow$ ViSQOL $\uparrow$ STOI $\uparrow$
Groundtruth 3.78 6.03 - -
@@
DM-Codec $\clubsuit$ 4.36 7.06 3.18 0.935
DM-Codec $\spadesuit$ 4.05 6.61 3.26 0.937
@@ 4 Experimental Results and Discussion
4.1 Comparison of Speech Tokenization models
@@
We compared the quality of DM-Codec’s discrete speech representations by reconstructing speech from quantized vector features and comparing it with state-of-the-art (SOTA) speech tokenization models: EnCodec, SpeechTokenizer, and FACodec. For LM-guided distillation, we utilize quantized features from the first Residual Vector Quantizer layer (RVQ-1), and for the combined LM and SM-guided distillation, we average all layers (RVQ-1:8).
Results: The results in Table 1 show that DM-Codec outperformed all evaluated SOTA speech tokenization models across most metrics. Specifically, DM-Codec with only LM-guided distillation exceeds the SOTA models, achieving improved scores: WER 4.36, WIL 7.06, and ViSQOL 3.18. Furthermore, DM-Codec’s with combined LM and SM-guided distillation outscore LM-guided distillation and all previous scores with 4.05 WER, 6.61 WIL, 3.26 ViSQOL, and achieved highly compatible 0.937 STOI scores compared to SOTA models.
Discussion: The observed performance gains stem from the proposed LM-guided distillation, which enhances the quantized features by leveraging LM’s contextual representations. This process aligns the speech with its overall context and word relation, resulting in more accurate reconstructions, as reflected in the reduced WER and WIL scores. By embedding contextual cues, the method effectively grounds isolated phonetic units within their overall context, reconstructing speech that aligns with human expectations, as demonstrated by the higher ViSQOL and STOI scores.
Moreover, the integration of LM and SM-based distillation further amplifies these improvements. The addition of SM distillation contributes to enhanced semantic-acoustic fidelity, as SM models capture phonetic nuances alongside prosodic and tonal characteristics. This dual representation—context from LM and phonetic detail from SM—produces a more coherent and natural speech reconstruction, yielding superior results across all metrics.

### v2

@@ 3.2 DM-Codec: Model Details
Our framework builds on the Residual Vector Quantizer with Generative Adversarial Networks (RVQ-GAN) architecture. The core model comprises an Encoder $\mathbf{E}$ and Decoder $\mathbf{D}$ with an RVQ structure, inspired by Encodec (Défossez et al., 2022) and SpeechTokenizer (Zhang et al., 2024a). We utilize a codebook size of 1024 and 8 quantization levels at a 50Hz frame rate. We employ a multi-discriminator setup, including Multi-Scale Discriminator (MSD), Multi-Period Discriminator (MPD), and Multi-Scale Short-Time Fourier Transform (MS-STFT) Discriminator, drawn from HiFi-Codec (Yang et al., 2023) and HiFi-GAN (Kong et al., 2020). Detailed architectural specifications are in Appendix E. To enhance the quantizer with distilled multimodal representations, we use wav2vec 2.0 (wav2vec2-base-960h) as $M_{ASR}$ (Baevski et al., 2020), BERT (bert-base-uncased) as $M_{LM}$ (Devlin et al., 2019), and HuBERT (hubert-base-ls960) as $M_{SM}$ (Hsu et al., 2021). Quantized outputs from the first RVQ layer (RVQ-1) are used for LM-guided distillation, and the average of quantized outputs across all eight layers (RVQ-1:8) is used for SM-guided distillation. An ablation study of RVQ layers selection is in Appendix D.2.
3.3 DM-Codec: Training Objective
@@
In addition to the distillation losses in Section 3.1, our training strategy builds on methodologies (Zhang et al., 2024a; Yang et al., 2023), employing the RVQ-GAN framework. For the original speech $\mathbf{x}$ and the reconstructed speech $\hat{\mathbf{x}}$, we use reconstruction, adversarial, feature matching, and commitment losses to guide learning as follows:
@@
Reconstruction Loss. To ensure that the model preserves the key attributes of speech, we employ both time-domain and frequency-domain reconstruction losses. The time-domain loss $\mathcal{L}_{t}$ is computed as the L1 distance between $\mathbf{x}$ and $\hat{\mathbf{x}}$. For the frequency-domain loss $\mathcal{L}_{f}$, we combine L1 and L2 losses over 64-bin Mel-spectrograms $\text{Mel}_{i}$, with varying window sizes of $2^{i}$, hop lengths of $2^{i}/4$, and scales $e=\{5,\ldots,11\}$.
@@
$\mathcal{L}_{f}$ $=\sum_{i\in e}\Big(\|\text{Mel}_{i}(\mathbf{x})-\text{Mel}_{i}(\hat{\mathbf{x}})\|_{1}$ (5)
$\quad+\|\text{Mel}_{i}(\mathbf{x})-\text{Mel}_{i}(\hat{\mathbf{x}})\|_{2}\Big)$
Adversarial Loss. The adversarial loss encourages the generator to produce realistic, indistinguishable speech. We apply a hinge loss formulation to compute the adversarial loss for the generator $\mathcal{L}_{g}$ and the discriminator $\mathcal{L}_{d}$. These losses are computed across all three discriminators: the multi-scale discriminator (MSD), multi-period discriminator (MPD), and the multi-scale STFT guided (MS-STFT) discriminator (see Appendix E).
@@
$\mathcal{L}_{d}$ $=\frac{1}{N}\sum_{n=1}^{N}\Big(\max(1-R_{n}(\mathbf{x}),0)$ (7)
$\quad+\max(1+R_{n}(\hat{\mathbf{x}}),0)\Big)$
@@
$\mathcal{L}_{G}$ $=\lambda_{L/LS}\mathcal{L}_{L/LS}+\lambda_{t}\mathcal{L}_{t}+\lambda_{f}\mathcal{L}_{f}$ (10)
$\quad+\lambda_{g}\mathcal{L}_{g}+\lambda_{fm}\mathcal{L}_{fm}+\lambda_{w}\mathcal{L}_{w}$
@@
3.4 DM-Codec-TTS: Model Details and Training Objective
Following SpeechTokenizer (Zhang et al., 2024a) and VALL-E (Wang et al., 2023), we propose DM-Codec-TTS, a novel multimodal representation distilled neural codec-based Text-To-Speech (TTS) model. Extending upon general neural codec language models, DM-Codec-TTS, leverages the strength of contextual and semantic representation distilled on our neural codec, DM-Codec.
Problem Formulation. For zero-shot TTS, the task is to synthesize speech for a given speaker. We frame it as a conditional codec language modeling problem, where the objective of DM-Codec-TTS is to predict the quantized acoustic features $\mathbf{Q}=\{\mathbf{Q}_{k}\}_{k=1}^{K}$, conditioned on a phoneme sequence $\mathbf{u}$ and an acoustic prompt $\tilde{\mathbf{P}}\in\mathbb{R}^{T^{\prime}\times K}$ extracted from the input enrolled recording.
Training Objective. The model integrates autoregressive (AR) and non-autoregressive (NAR) components to hierarchically encode speech information. The AR component models content and speaker identity by predicting tokens $\mathbf{Q}_{1}^{t}$ from the first RVQ layer using a transformer decoder-only architecture $\phi_{\text{AR}}$. The AR training objective is:
$\mathcal{J}_{\text{AR}}=-\sum_{t=0}^{T}\log\,p\left(\mathbf{Q}_{1}^{t}\mid\mathbf{Q}_{1}^{<t},\mathbf{u};\phi_{\text{AR}}\right)$ (11)
The NAR component focuses on acoustic details by predicting tokens $\mathbf{Q}_{k}$ ($k=2,\dots,8$) from subsequent RVQ layers. The NAR training objective is:
$\mathcal{J}_{\text{NAR}}=-\sum_{k=2}^{8}\log\,p\left(\mathbf{Q}_{k}\mid\mathbf{Q}_{<k},\tilde{\mathbf{P}},\mathbf{u};\phi_{\text{NAR}}\right)$ (12)
During inference, the AR model predicts tokens $\mathbf{Q}_{1}$ based on speech phonemes $\mathbf{u}$, while the NAR model iteratively generates $\mathbf{Q}_{2:8}$ using the AR output and the acoustic prompt $\tilde{\mathbf{P}}$. The combined tokens $\mathbf{Q}=\{\mathbf{Q}_{1},\mathbf{Q}_{2:8}\}$ are decoded into a speech waveform via the neural codec.
Model Details. The AR and NAR models share an identical transformer structure, comprising 12 layers, 16 attention heads, a 1024-dimensional embedding space, 4096-dimensional feed-forward layers, and a dropout rate of 0.1.
@@ 3.4 DM-Codec-TTS: Model Details and Training Objective
Model WER $\downarrow$ WIL $\downarrow$ ViSQOL $\uparrow$ STOI $\uparrow$ Similarity $\uparrow$ MOS $\uparrow$ UTMOS $\uparrow$ PESQ $\uparrow$
Groundtruth 3.78 6.03 - - - - - -
@@
[…]
DM-Codec (Baseline) 4.97 8.02 2.95 0.935 0.991 3.13 3.25 2.58
DM-Codec (SM Baseline) 4.49 7.25 3.12 0.933 0.994 - 3.43 2.86
DM-Codec (CLS) 4.47 7.08 3.12 0.926 0.993 3.65 3.39 2.65
DM-Codec (LM) 4.36 7.06 3.18 0.935 0.994 3.69 3.48 2.86
DM-Codec (LM+SM) 4.05 6.61 3.26 0.937 0.994 3.72 3.52 2.81
4 Experimental Setup
Dataset. We trained DM-Codec using the LibriSpeech 100-hour clean speech training set (Panayotov et al., 2015), widely used for speech tokenizer and modeling (Zhang et al., 2024a; Ju et al., 2024; Hsu et al., 2021). Data were standardized by cropping samples to three seconds and ensuring a 16 Hz sample rate. For DM-Codec-TTS, we used LibriHeavy (Kang et al., 2024), a 50k-hour read English speech dataset derived from LibriVox, selecting samples between 0.5 and 15 seconds at 16 kHz. Additionally, a smaller DM-Codec-TTS version was trained on LibriTTS (Zen et al., 2019), a 585-hour dataset with 2,456 speakers sampled at 24 kHz.
For evaluation, we tested DM-Codec on 300 randomly selected audio samples from the LibriSpeech test-clean subset, aligning with baseline Speechtokenizer’s (Zhang et al., 2024a) setup. A random seed of 42 ensured consistent sampling, and baseline models were evaluated on the same setup. Additionally, we tested DM-Codec-TTS on LibriSpeech test-clean and VCTK (Yamagishi et al., 2019). For LibriSpeech, we followed VALL-E’s (Wang et al., 2023) setup by constructing a 2.2-hour subset from the test-clean data, selecting samples of 4–10 seconds. For each synthesis, a different utterance from the same speaker was randomly selected as input text, and a 3-second segment was cropped as the enrollment speech prompts. Each experiment was repeated three times, and the average score was reported. For VCTK, we followed SpeechTokenizer’s (Zhang et al., 2024a) setup, using a 3s utterance as the prompts, while the input text was derived from a different utterance.
Training. We trained DM-Codec utilizing two A100 GPUs for 100 epochs with batch size of 6. We applied a learning rate of $1\times 10^{-4}$ using the Adam optimizer with a 0.98 learning rate decay. The embedding size was set to 1024 for RVQ and 768 for the LM and SM. For all experiments, we used a random seed of 42 to ensure reproducibility. For the overall generator loss, we select the weight coefficients proportionally as follows: $\lambda_{L/LS}=X$, $\lambda_{f}=0.375X$, $\lambda_{t}=4.15X$, $\lambda_{w}=0.085X$, and set $\lambda_{g}=1$ and $\lambda_{fm}=1$. For DM-Codec-TTS, we trained the autoregressive (AR) and non-autoregressive (NAR) models separately, each for 4 epochs, while DM-Codec-TTS-small was trained with 100 epochs for the AR model and 196 epochs for the NAR model. The batch size was determined dynamically based on the maximum number of audio seconds, set to 280 for AR and 200 for NAR. We employed a base learning rate of 0.05 using the ScaledAdam optimizer with warmup steps of 200.
Baselines. We compared DM-Codec with baselines: EnCodec (24 kHz) (Défossez et al., 2022), SpeechTokenizer (Zhang et al., 2024a), FACodec (NaturalSpeech3) (Ju et al., 2024), DAC (16 kHz) (Kumar et al., 2023), BigCodec (Xin et al., 2024), and Moshi (Défossez et al., 2024). We reproduced SpeechTokenizer using its official training code and used official model checkpoints for the rest. Additionally, we compared DM-Codec-TTS with neural codec language models: USLM (from SpeechTokenizer (Zhang et al., 2024a)) and VALL-E (Wang et al., 2023). Since VALL-E and USLM’s training codes and models are not open-source, we relied on reported results. We also used USLM (libri), an official model checkpoint trained on LibriTTS and shared on GitHub. All baseline model artifacts are open source and were used according to their licenses and intended research use.
Evaluation Metrics. To measure context preservation, we used Word Error Rate (WER) and Word Information Lost (WIL), with Whisper (whisper-medium) (Radford et al., 2023) generating the transcriptions. WER quantifies transcription errors, while WIL measures the loss of key information:
$\text{WER}=\frac{S+D+I}{N},\quad\text{WIL}=1-\frac{N-S-D}{N}$
where $N$ is the total number of words in the reference, $S$ is the number of correctly recognized words, $D$ is the number of deletions, and $I$ is the number of insertions. Groundtruth WER and WIL scores were included to account for Whisper’s transcription errors. We assess the acoustic and semantic preservation using ViSQOL (Virtual Speech Quality Objective Listener) (Hines et al., 2012) measuring spectro-temporal similarity, Short-Time Objective Intelligibility (STOI) measuring short-time correlations, and and Perceptual Evaluation of Speech Quality (PESQ) measuring speech quality based on perceptual auditory models. We measured the Mean Opinion Score (MOS) and Similarity Mean Opinion Score (SMOS) through human evaluations with 50 English-proficient participants. Evaluators rated anonymized samples on a 1-to-5 scale. MOS is used to evaluate the naturalness, intelligibility, and clarity of speech, while SMOS is used to measure the similarity to the prompt speaker’s voice.We calculated UTMOS (Saeki et al., 2022), a neural network-based automatic MOS predictor to further complement subjective MOS. Additionally, speaker similarity (Similarity) was quantified using cosine similarity between normalized speaker embeddings extracted with WavLM-TDNN (Chen et al., 2022). Further details on human evaluations are in Appendix F.
@@ 5 Experimental Results and Discussion
In this section, we present experimental results evaluating speech reconstruction with DM-Codec variants: DM-Codec (LM) with LM-guided distillation, DM-Codec (LM+SM) with combined LM and SM distillation, and DM-Codec (CLS) with [CLS] token guided distillation against baseline speech tokenizers (§5.1). We then assess zero-shot speech synthesis with DM-Codec-TTS against neural codec language models (§5.2). Furthermore, we conducted rigorous significance analysis (Appendix C) and comprehensive ablation studies of DM-Codec design choices (Appendix D).
Table 2: Evaluation of DM-Codec-TTS on LibriSpeech and VCTK datasets. Results show DM-Codec-TTS outperforms baselines in WER, WIL, Similarity, MOS, and SMOS. $\diamondsuit$ denotes results from official model checkpoints, $\dagger$ indicates results from the paper, and $\clubsuit$ indicates models trained with LibriTTS dataset. Bold highlights best result.
LibriSpeech Benchmark Evaluation VCTK Benchmark Evaluation
Model WER $\downarrow$ WIL $\downarrow$ Similarity $\uparrow$ MOS $\uparrow$ SMOS $\uparrow$ WER $\downarrow$ WIL $\downarrow$ Similarity $\uparrow$ MOS $\uparrow$ SMOS $\uparrow$
DM-Codec-TTS 5.08 7.32 0.82 3.70 3.89 3.58 5.65 0.82 3.78 3.85
Vall-E $\dagger$ 5.90 - 0.58 - 4.38 - - 0.38 - 3.81
USLM $\dagger$ - - - - 6.50 - 0.84 3.63 3.45
@@
[…]
USLM (libri) $\diamondsuit$ $\clubsuit$ 16.72 25.65 0.80 3.11 2.83 14.79 23.24 0.78 2.94 2.63
5.1 Speech Reconstruction Evaluation
@@
We compared the quality of DM-Codec’s discrete speech representations by reconstructing speech from quantized features. For this evaluation, we selected DM-Codec variants with the first Residual Vector Quantizer layer (RVQ-1) for LM distillation and all RVQ layers (RVQ-1:8) for SM distillation. We analyze the results of DM-Codec in effectively incorporating semantic and contextual information.
Results: Table 1 shows that all DM-Codec variants surpass or closely compete with the baselines. DM-Codec (LM+SM) outperforms all models in reducing transcription error (WER 4.05, WIL 6.61) and speech quality (ViSQOL 3.26, MOS 3.72, UTMOS 3.52). DM-Codec (LM) surpasses SpeechTokenizer, EnCodec, BigCodec, FACodec, and Mimi in reducing transcription error (WER 4.36, WIL 7.06) and speech quality (ViSQOL 3.18, PESQ 2.86), while achieving a higher UTMOS (3.69) than DAC, EnCodec, and Mimi. Additionaly, DM-Codec (CLS) demonstrates commendable results, either outperforming or competing with the baselines, particularly outscoring SpeechTokenizer, EnCodec, BigCodec, FACodec, and Mimi in reducing transcription error (WER 4.47, WIL 7.08) and SpeechTokenizer, EnCodec, and Mimi in speech quality (ViSQOL 3.12, PESQ 2.65), while achieving a strong MOS of 3.65.
Discussion: The superior performance of DM-Codec (LM+SM) is attributed to the novel Combined LM and SM-guided distillation. This dual representation, contextual knowledge from LM and semantic understanding from SM enables more coherent and natural speech, yielding superior results across most evaluation metrics. DM-Codec (LM) also performs well by incorporating contextual representations, improving speech quality and reducing transcription errors. Additionally, [CLS] token-guided distillation provides a holistic view of the entire contextual input, enhancing contextual cue alignment and reducing transcription errors. The impact of these distillation techniques is clear when compared to DM-Codec (Baseline), which lacks distillation and falls short in speech quality with higher transcription errors. Moreover, the importance of contextual information is reinforced by the comparison with DM-Codec (SM Baseline), which only includes semantic knowledge and underperforms with more transcription errors.
5.2 Speech Synthesis Evaluation
To demonstrate DM-Codec’s ability to capture semantic and contextual features and their impact on downstream task, we compare the zero-shot TTS evaluation of DM-Codec-TTS with baseline USLM and VALL-E. We use DM-Codec (LM+SM) to quantize speech prompt features as input and decode quantized features predicted by DM-Codec-TTS. DM-Codec(SM)-TTS serves as a baseline.
Results: The results in Table 2 show that DM-Codec-TTS outperforms USLM and VALL-E baselines. In both benchmark evaluations, DM-Codec achieves the lowest WER (5.08, 3.58), WIL (7.32, 5.65), and highest MOS (3.70, 3.78), while achieving closely aligned Similarity to USLM in VCTK, and superior SMOS 3.85 compared to VALL-E and USLM in VCTK. DM-Codec-TTS trained with smaller LibriTTS dataset also significantly outperforms USLM (libri) in both benchmarks.
Discussion: The improved performance indicates that hierarchical modeling in DM-Codec-TTS effectively utilizes the contextual and semantic knowledge distilled in DM-Codec. Unlike VALL-E, which relies on EnCodec for speech tokenization, DM-Codec-TTS leverages multimodal representation cues. This highlights the strength of contextual and semantic-aware hierarchical modeling in bridging linguistic content with acoustic fidelity, resulting in more natural and intelligible speech synthesis. The improvement over the DM-Codec(SM)-TTS baseline also demonstrates the importance of contextual distillation from LM.
6 Conclusion
We introduced DM-Codec, a speech tokenizer with novel distillation methods that leverage multimodal (acoustic, semantic, and contextual) representations from language and speech self-supervised models. Experimental results demonstrate that distilling multimodal representations enables DM-Codec to encode salient speech information in discrete tokens. This approach showcases the potential of multimodal representations to improve speech tokenization across domains, including multilingual and code-switched speech processing.
Limitations
In this work, we present the effectiveness of our proposed method, DM-Codec, based on the LibriSpeech dataset. Future research could investigate its performance across a variety of datasets and domains. Additionally, exploring the capabilities of DM-Codec in multilingual contexts would be valuable. Another limitation of our work is the absence of experiments with emerging LLMs. Currently, we focus solely on masked language models to derive representations. Further investigation into decoder-based LLMs’ impact on DM-Codec can be studied and addressed.
Acknowledgments
This project was partially supported by a grant from Independent University, Bangladesh (IUB).
Appendix A Resources
We release all resources to facilitate reproducibility and future research. We provide the code for training DM-Codec, trained model checkpoints for inference, and a Dockerfile for a reproducible environment. All resources are publicly available at https://github.com/mubtasimahasan/DM-Codec.